### v3

@@ 1 Introduction
Recent advancements in deep learning, particularly in Convolutional Neural Networks (CNNs), have shown significant promise in improving pneumonia diagnosis from chest X-rays yan2023pneumonia ; lin2023enhancing . These models can analyze medical images with remarkable precision, often outperforming human radiologists in both consistency and speed. Recent innovations, such as attention mechanisms, have further enhanced diagnostic accuracy askari2024enhancing ; parhami2023comparison ; fateh2024advancing . Additionally, transformers have demonstrated significant potential in medical imaging tasks due to their ability to model long-range dependencies and identify complex patterns aurangzeb2024introducing ; fateh2024msdnet . These advancements highlight the potential of AI to complement radiologists’ expertise and enhance patient outcomes.
Lung segmentation is a critical preprocessing step for improving the accuracy of pneumonia detection in chest X-rays. However, this task faces several challenges. First, the presence of artifacts, overlapping anatomical structures, and low contrast in chest X-rays can make it difficult to accurately delineate lung boundaries alam2024attention . Traditional segmentation models like U-Net, while effective, often struggle to capture fine-grained details and contextual information, leading to suboptimal performance in complex cases zhang2023cams . Additionally, the high variability in lung shapes and sizes across patients further complicates the segmentation process. Although transformer-based models have shown promise in addressing these issues, they often come with high computational costs and large parameter counts, making them unsuitable for resource-constrained environments cao2022swin .
Classification of pneumonia from chest X-rays also presents significant challenges. First, the subtle visual indicators of pneumonia, such as small opacities or localized consolidations, can be easily missed by traditional CNN-based models like ResNet and DenseNet, which primarily focus on local features rajpurkar2017chexnet . While transformer-based models excel at capturing global contextual information, they require large amounts of labeled data and extensive computational resources, limiting their practicality in real-world clinical settings rezvani2024fusionlungnet . Moreover, the domain shift between pre-trained models (e.g., on ImageNet) and medical imaging datasets often results in suboptimal performance, necessitating advanced techniques like transfer learning and domain adaptation chen2017rethinking . These challenges highlight the need for a computationally efficient and accurate model that can effectively leverage both local and global features for pneumonia diagnosis.
@@ 2.1.1 U-Net for CXR Segmentation
The U-Net architecture, with its encoder-decoder structure and skip connections, has occurred as a leading method for CXR segmentation. This setup, which captures high-level semantic information and low-level details, is crucial for accurately outlining lung boundaries. Studies have consistently shown U-Net’s effectiveness in segmenting lung regions with high accuracy, a factor that significantly comforts the potential of this technology in improving diagnostic outcomes fateh2023persian ; fateh2022providing . U-Net, introduced by Ronneberger et al., has become a fundamental tool in medical image segmentation ronneberger2015u . Islam et al. islam2018towards showcased U-Net’s ability in accurately tracing lung boundaries, which has enhanced diagnostic precision. Additionally, Liu et al. liu2022automatic employed a pre-trained EfficientNet-B4 and developed an enhanced version of U-Net for identifying and segmenting lung regions.
@@ 2.1.2 U-Net Enhancements with Transformers
To address traditional U-Net limitations, advanced architectures that enhance U-Net’s ability to capture multi-scale and contextual information are needed. Recent research has significantly advanced lung segmentation by enhancing the U-Net architecture with attention mechanisms. Oktay et al. oktay2018attention introduced mechanisms that enable the model to concentrate on the most crucial areas within chest X-rays using Attention Gates (AGs). This innovation enhances segmentation accuracy and sensitivity to disease characteristics. Additionally, research by Wu et al. wu2023transformer , Gu et al. gu2018recent , and Liu et al. liu2024sta has shown that incorporating attention mechanisms into the U-Net framework significantly improves lung segmentation performance, underscoring the effectiveness of this method in precisely outlining lung boundaries and enhancing diagnostic outcomes.
Khaniki et al. khaniki2024novel enhanced U-Net by incorporating a Convolutional Block Attention Module (CBAM), which integrates channel, spatial, and pixel attention to boost segmentation accuracy. Azad et al. and Chen et al. extended the U-Net framework with transformers, demonstrating significant improvements in capturing intricate details and achieving top-tier results in lung segmentation tasks azad2019bi ; chen2021transunet .
The incorporation of transformer modules has marked a landmark in lung segmentation research. Transformer architectures, known for capturing long-range dependencies and contextual information from text, have been successfully integrated into U-Net variants, leading to notable improvements in segmentation accuracy. For instance, Chen et al. chen2023cotrfuse created a hybrid CNN-Transformer model for medical image segmentation, merging the strengths of CNNs and transformers to enhance accuracy and robustness in lung tissue segmentation. Similarly, Valanarasu et al. valanarasu2021medical developed hybrid models combining U-Net with transformer modules, effectively utilizing CNNs and transformers to capture complex anatomical details and spatial relationships in chest X-ray images.
@@ 2.2.1 Classical Approaches for CXR Classification
Early methods for classifying chest X-ray (CXR) images primarily depended on traditional machine learning techniques, employing classifiers such as Support Vector Machines (SVM), K-nearest Neighbors (k-NN), and Random Forests. For example, Stokes et al. used logistic regression, decision trees, and SVM to categorize patients’ clinical data into bronchitis or pneumonia, with decision trees yielding the highest recall value of 80% and an AUC of 93% stokes2021machine . Similarly, Qi et al. applied logistic regression and random forest to CT images, achieving AUCs of 0.97 and 0.92, respectively qi2020machine . Chandra et al. used a multi-layer perceptron (MLP) to segment lung regions from CXR images, reaching an accuracy of 95.39% chandra2020pneumonia . However, these methods, which heavily relied on symptomatic data, had limited accuracy and were evaluated on small datasets wang2024prediction ; berg2023employing ; fateh2021multilingual .
@@ 2.2.2 Deep Learning Models
The beginning of deep learning, especially Convolutional Neural Networks (CNNs), has significantly transformed medical image analysis by providing superior accuracy and robustness sanghvi2023deep ; shiri2024differentiation rajinikanth2022pneumonia ; allioui2022multi . For instance, Stephen et al. designed a custom CNN model from scratch, achieving a training accuracy of 95.31% and a validation accuracy of 93.73% stephen2019efficient . Similarly, Sharma et al. created a straightforward CNN architecture that reached a 90.68% accuracy rate on the ”Kermany” dataset using data augmentation sharma2020feature ; kermany2018labeled . Rehman et al rehman2021self , used self-activated neural networks for multi-class classification to detect chest diagnosis. However, relying solely on data augmentation does not introduce substantially new information, restricting the model’s ability to learn a wide range of complex patterns from the training data.
@@ 2.2.3 Transfer Learning
Pre-trained CNNs have become the standard for image classification tasks, including CXR analysis. These models leverage large datasets and transfer learning to enhance performance on specific medical imaging tasks. Transfer learning, where pre-trained models are adapted and refined for new, specific tasks, has achieved significant results. For instance, Rajpurkar et al. utilized DenseNet-121 on the ChestX-ray8 dataset, comprising 112,150 frontal CXR images, achieving an F1-score of 76.8%. This study highlighted the potential of transfer learning in medical image classification rajpurkar2017chexnet . Choudhary et al. used customized VGG19 and transfer learning to classify Optical Coherence Tomography (OCT), achieving 99.17% accuracy choudhary2023deep . Rahman et al. rahman2020transfer utilized transfer learning techniques exclusively, employing various CNN models pre-trained on ImageNet data to classify pneumonia.
@@ 2.2.4 Ensemble Approaches
Ensemble learning, which combines the outputs of multiple CNN models, has shown considerable promise. For instance, Ukwuoma et al. ukwuoma2023hybrid proposed two ensemble methods: ensemble group A (DenseNet201, VGG16, and GoogleNet) and ensemble group B (DenseNet201, InceptionResNetV2, and Xception). These models, followed by a self-attention layer and a multi-layer perceptron (MLP) for disease identification, achieved 97.22% accuracy for binary classification, and 97.2% and 96.4% for multi-class classification, respectively. Kundu et al. kundu2021pneumonia combined GoogLeNet, ResNet-18, and DenseNet-121 in an ensemble model, reaching an accuracy of 86.85%, albeit with high computational costs. Jaiswal et al. jaiswal2019identifying used a mask region-based CNN for pneumonia detection through segmentation, employing an ensemble of ResNet-50 and ResNet-101 for image thresholding. In the RSNA Pneumonia Detection Challenge, Gabruseva et al. gabruseva2020deep presented a deep learning framework using the single-shot detector RetinaNet with Se-ResNext101 encoders, achieving a mean average precision (mAP) of 0.26 through snapshot ensembling.
@@ 2.2.5 Transformers
Recent advancements in medical image classification have harnessed transformer architectures alongside deep learning, yielding impressive outcomes. Wang et al. wang2021transpath unveiled TransPath, a hybrid model merging CNN and transformer architectures, highlighting the potential of such integrations. They proved the efficacy of self-supervised pretraining on extensive datasets like TCGA and PAIP, followed by fine-tuning on specific medical image datasets, resulting in solid performance: 89.68% accuracy on MHIST, 95.85% on NCT-CRC-HE, and 89.91% on PatchCamelyon. Transformer-based models have garnered attention for their capacity to capture long-range dependencies in images. Wu et al. wu2022swin introduced a Swin Transformer-based model for pulmonary nodule classification, successfully adapting the architecture to the smaller scale of medical image datasets and achieving significant results. Dai et al. dai2021transmed investigated the use of transformers for multimodal medical image classification with their TransMed model, reaching 88.9% accuracy on the PGT dataset and 85% on the MRNet dataset. Leamons et al. leamons2022vision concentrated on breast cancer detection, comparing CNNs, RNNs, and Visual Transformers (VTs), and found that the VT model excelled with a 93% accuracy. Jang et al. jang2022m3t proposed M3T, a 3D medical image classifier that combines transformers with 2D and 3D CNNs, showcasing its effectiveness across various medical datasets.
@@
Hybrid architectures that combine CNNs and transformers have demonstrated remarkable success in pneumonia detection. For instance, a hybrid model integrating ResNet34 with a Multi-Axis Vision Transformer achieved a state-of-the-art accuracy of 94.87% on the Kaggle pediatric pneumonia dataset. This model leverages the local feature extraction capabilities of CNNs and the global context modeling of transformers, resulting in fewer misclassifications and improved robustness angara2024novel . Additionally, another study proposed a hybrid CNN-Swin Transformer model that achieved 98.72% accuracy, demonstrating the effectiveness of combining convolutional layers with transformer blocks for enhanced feature extraction mustapha2025enhanced .
@@ 4.4 Comparison with State-of-the-Art
FusionLungNet rezvani2024fusionlungnet 95.29% 97.73% 98.66% 92.14% 95.29% 93.89%

### v4

@@ 1 Introduction
Recent advancements in deep learning, particularly in Convolutional Neural Networks (CNNs), have shown significant promise in improving pneumonia diagnosis from chest X-rays lin2023enhancing . These models can analyze medical images with remarkable precision, often outperforming human radiologists in both consistency and speed. Recent innovations, such as attention mechanisms, have further enhanced diagnostic accuracy askari2025enhancing ; parhami2023comparison ; fateh2024advancing . Additionally, transformers have demonstrated significant potential in medical imaging tasks due to their ability to model long-range dependencies and identify complex patterns fateh2024msdnet . These advancements highlight the potential of AI to complement radiologists’ expertise and enhance patient outcomes.
Lung segmentation is a critical preprocessing step for improving the accuracy of pneumonia detection in chest X-rays. However, this task faces several challenges. First, the presence of artifacts, overlapping anatomical structures, and low contrast in chest X-rays can make it difficult to accurately delineate lung boundaries sharifi2024da ; agnes2024wavelet . Traditional segmentation models like U-Net, while effective, often struggle to capture fine-grained details and contextual information, leading to suboptimal performance in complex cases yang2024novel . Additionally, the high variability in lung shapes and sizes across patients further complicates the segmentation process. Although transformer-based models have shown promise in addressing these issues, they often come with high computational costs and large parameter counts, making them unsuitable for resource-constrained environments narmadha2025het .
Classification of pneumonia from chest X-rays also presents significant challenges. First, the subtle visual indicators of pneumonia, such as small opacities or localized consolidations, can be easily missed by traditional CNN-based models like ResNet and DenseNet, which primarily focus on local features amin2024automated . While transformer-based models excel at capturing global contextual information, they require large amounts of labeled data and extensive computational resources, limiting their practicality in real-world clinical settings rezvani2025fusionlungnet . Moreover, the domain shift between pre-trained models (e.g., on ImageNet) and medical imaging datasets often results in suboptimal performance, necessitating advanced techniques like transfer learning and domain adaptation chen2017rethinking . These challenges highlight the need for a computationally efficient and accurate model that can effectively leverage both local and global features for pneumonia diagnosis.
@@ 2.1.1 U-Net for CXR Segmentation
The U-Net architecture, with its encoder-decoder structure and skip connections, has occurred as a leading method for CXR segmentation. This setup, which captures high-level semantic information and low-level details, is crucial for accurately outlining lung boundaries. Studies have consistently shown U-Net’s effectiveness in segmenting lung regions with high accuracy, a factor that significantly comforts the potential of this technology in improving diagnostic outcomes fateh2023persian ; fateh2022providing . U-Net, introduced by Ronneberger et al., has become a fundamental tool in medical image segmentation ronneberger2015u . Additionally, Liu et al. liu2022automatic employed a pre-trained EfficientNet-B4 and developed an enhanced version of U-Net for identifying and segmenting lung regions.
@@ 2.1.2 U-Net Enhancements with Transformers
To address traditional U-Net limitations, advanced architectures that enhance U-Net’s ability to capture multi-scale and contextual information are needed. Recent research has significantly advanced lung segmentation by enhancing the U-Net architecture with attention mechanisms. Oktay et al. oktay2018attention introduced mechanisms that enable the model to concentrate on the most crucial areas within chest X-rays using Attention Gates (AGs). This innovation enhances segmentation accuracy and sensitivity to disease characteristics.
Azad et al. and Chen et al. extended the U-Net framework with transformers, demonstrating significant improvements in capturing intricate details and achieving top-tier results in lung segmentation tasks chen2021transunet .
The incorporation of transformer modules has marked a landmark in lung segmentation research. Transformer architectures, known for capturing long-range dependencies and contextual information from text, have been successfully integrated into U-Net variants, leading to notable improvements in segmentation accuracy. For instance, Chen et al. chen2023cotrfuse created a hybrid CNN-Transformer model for medical image segmentation, merging the strengths of CNNs and transformers to enhance accuracy and robustness in lung tissue segmentation.
@@ 2.2.1 Classical Approaches for CXR Classification
Early methods for classifying chest X-ray (CXR) images primarily depended on traditional machine learning techniques, employing classifiers such as Support Vector Machines (SVM), K-nearest Neighbors (k-NN), and Random Forests. For example, Stokes et al. used logistic regression, decision trees, and SVM to categorize patients’ clinical data into bronchitis or pneumonia, with decision trees yielding the highest recall value of 80% and an AUC of 93% stokes2021machine . Chandra et al. used a multi-layer perceptron (MLP) to segment lung regions from CXR images, reaching an accuracy of 95.39% chandra2020pneumonia . However, these methods, which heavily relied on symptomatic data, had limited accuracy and were evaluated on small datasets wang2024prediction ; fateh2021multilingual .
@@ 2.2.2 Deep Learning Models
The beginning of deep learning, especially Convolutional Neural Networks (CNNs), has significantly transformed medical image analysis by providing superior accuracy and robustness allioui2022multi . For instance, Stephen et al. designed a custom CNN model from scratch, achieving a training accuracy of 95.31% and a validation accuracy of 93.73% stephen2019efficient . Similarly, Sharma et al. created a straightforward CNN architecture that reached a 90.68% accuracy rate on the ”Kermany” dataset using data augmentation kermany2018labeled . However, relying solely on data augmentation does not introduce substantially new information, restricting the model’s ability to learn a wide range of complex patterns from the training data.
@@ 2.2.3 Transfer Learning
Pre-trained CNNs have become the standard for image classification tasks, including CXR analysis. These models leverage large datasets and transfer learning to enhance performance on specific medical imaging tasks. Transfer learning, where pre-trained models are adapted and refined for new, specific tasks, has achieved significant results. For instance, Rajpurkar et al. utilized DenseNet-121 on the ChestX-ray8 dataset, comprising 112,150 frontal CXR images, achieving an F1-score of 76.8%. This study highlighted the potential of transfer learning in medical image classification rajpurkar2017chexnet .
@@ 2.2.4 Ensemble Approaches
Ensemble learning, which combines the outputs of multiple CNN models, has shown considerable promise. For instance, Ukwuoma et al. ukwuoma2023hybrid proposed two ensemble methods: ensemble group A (DenseNet201, VGG16, and GoogleNet) and ensemble group B (DenseNet201, InceptionResNetV2, and Xception). These models, followed by a self-attention layer and a multi-layer perceptron (MLP) for disease identification, achieved 97.22% accuracy for binary classification, and 97.2% and 96.4% for multi-class classification, respectively. Jaiswal et al. jaiswal2019identifying used a mask region-based CNN for pneumonia detection through segmentation, employing an ensemble of ResNet-50 and ResNet-101 for image thresholding.
@@ 2.2.5 Transformers
Recent advancements in medical image classification have harnessed transformer architectures alongside deep learning, yielding impressive outcomes. Wang et al. wang2021transpath unveiled TransPath, a hybrid model merging CNN and transformer architectures, highlighting the potential of such integrations. They proved the efficacy of self-supervised pretraining on extensive datasets like TCGA and PAIP, followed by fine-tuning on specific medical image datasets, resulting in solid performance: 89.68% accuracy on MHIST, 95.85% on NCT-CRC-HE, and 89.91% on PatchCamelyon. Transformer-based models have garnered attention for their capacity to capture long-range dependencies in images. Wu et al. wu2022swin introduced a Swin Transformer-based model for pulmonary nodule classification, successfully adapting the architecture to the smaller scale of medical image datasets and achieving significant results.
@@
Hybrid architectures that combine CNNs and transformers have demonstrated remarkable success in pneumonia detection. For instance, a hybrid model integrating ResNet34 with a Multi-Axis Vision Transformer achieved a state-of-the-art accuracy of 94.87% on the Kaggle pediatric pneumonia dataset. This model leverages the local feature extraction capabilities of CNNs and the global context modeling of transformers, resulting in fewer misclassifications and improved robustness angara2024novel .
@@ 4.4 Comparison with State-of-the-Art
FusionLungNet rezvani2025fusionlungnet 95.29% 97.73% 98.66% 92.14% 95.29% 93.89%